Could you please help me solve this?:
The Ed chairlift has a vertical rise of 900 feet and the Green Machine chairlift has a vertical rise of 1200 feet. Both ski lifts start at the ski lodge and Jason started and ended his day at the ski lodge. Jason skied a total of 15,900 vertical feet and rode only these two chairlifts a total of 16 times. How many times did he ride each chairlift?

Answer by stanbon(48551)

The Ed chairlift has a vertical rise of 900 feet and the Green Machine chairlift has a vertical rise of 1200 feet. Both ski lifts start at the ski lodge and Jason started and ended his day at the ski lodge. Jason skied a total of 15,900 vertical feet and rode only these two chairlifts a total of 16 times. How many times did he ride each chairlift?
-------------------
This is like a coin problem where the value of the two coins is
900 ft or 1200 ft.
----------
Let Jason's number of Ed runs be "x": Vert dist = 900x ft
Then Jason's number of Green runs is "16-x": Vert dist = 1200(16-x)
=19200-1200x ft
-----------------
EQUATION:
vert dist + vert dist = 15,900 vert ft.
900x + 19200-1200x = 15,900
-300x=-3300
x=11 (# of runs on Ed)
16-x=5 (# of runs on Green)
--------------
Cheers,
Stan H.